) John consumes two goods, X and Y. The marginal utility of X and the marginal utility of Y satisfy the following equations: MUX = Y MUY = X. The price of X is $9, and the price of Y is $12.

a. Write an expression for Johnʹs MRS.

b. What is the optimal mix between X and Y in Johnʹs market basket?

c. John is currently consuming 15 X and 10 Y per time period. Is he consuming an optimal mix of X and Y?
Answer #1

a. MRS = MUx / MUy

MRS = Y/X

b. At optimality condition

MRS= Px/ Py

Y/X = 9/12

Y= (3/4)X

Budget line

Px. X+ Py. Y= I

X= 4I/ (4.Px+ 3.Py)

Y= 3I/(4.Px+3.Py)

c. X=15 and Y=10

MRS= Y/X = 10/15 = 2/3 but price ratio is 3/4

Hence tangency condition is not met it is not optimal.